OpenStudy (anonymous):

A wave with polarization parallel to the plane of incident powers is transmitted from a medium with refractive index n1 to a medium with refractive index n2. We say that the Brewster angle in this case of 33.82 o. Which of the following statements is true regarding the critical angle? Justify the answers. (a) The critical angle is worth 31, 29o. (b) If there is Brewster angle, there can be no critical angle. (c) The critical angle is worth 42, 06o. (d) The critical angle it always π / 2.

OpenStudy (anonymous):

Definition of Brewster's angle is that the reflected light is prependicular to the transmited light.  So we have thita + phi = 90o. Then, by n1 sin thita = n2 sin phi = n2 sin (90-thita) = n2 cos thita, we get tan thita = tan 33.82o = n2/n1 = 0.67. The critical angle is arc-sin (n2/n1) = arc-sin (0.67) = 42.06o. I believe the answer is C